At Calcutta, here the vertical component of the earth's field is 0.26(dyne per unit pole) it is found that a mass of 10mg placed 10cm from the axel of a dip-needle keeps it horizontal.Find the magnetic moment of the dip-needle( acceleration due to gravity is 1000cm/sec2)

Grade:11

1 Answers

Askiitians Expert Soumyajit IIT-Kharagpur
13 years ago

Dear Sanket,

Ans:- Let the magnetic moment is M unit.

Then the moment about the axle of the dip needle is MV where V=0.26 dyne/pole

The moment given by the point mass=mgx where x=10 cm

Hence we get,

MV=mgx

M=mgx/V

  =(10*10-³ *10³ *10)/0.26=384.6 unit(ans)

 

Please feel free to post as many doubts on our discussion forum as you can. If you find any question
Difficult to understand - post it here and we will get you the answer and detailed solution very quickly. We
are all IITians and here to help you in your IIT JEE preparation.


All the best Sanket !!!

 


Regards,

Askiitians Experts

SOUMYAJIT IIT_KHARAGPUR